Hi! I hope to receive help with the following question and I thank you in advance! I will give you a thmbs up!

----------------------------

2) It is known that when the price of a good is £2, 50 units are demanded, but when the price increases to £4 the quantity demanded drops to 45 units (that is, a line contains the points (2, 50) and (4, 45)).

a) Calculate the change in demand that results from a 1 unit increase in price.

b) Calculate the quantity demanded at prices of £5, £8, and £10 per unit.

c) Graph the demand function.

d) Write down the equation of the demand function in the form,

(i) P = g (Q), and (ii) Q =f (P)
Answer #1

Please find the answer in the attached screenshot

ay ay - (2,50) (4,45) 94 Equation is ma 45-50 2-5 XX, 4-2 Z (4-50) = (3-2) 24-100 sato. sutry= 110 when x 25, 8, 102? 5(5) +

Demand function - c) Deman qrantiny 0 1 2 3 4 5 6 7 8 9 TOS price 24 z 110-50 5x+2y = 110. 502 110-24 dz 110-24 y z 10-50 Pzg